Grade 11Algebra

iz^3+z^2-z+i=0 ,then find out |z| value.plz ans. Of my question properly.

CA
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
iz^3 + z^2 - z + i = 0 
==> iz^2 (z - i) - 1(z - i) = 0, factoring by grouping 
==> (iz^2 - 1) (z - i) = 0 
==> i(z^2 + i) (z - i) = 0 
==> z^2 + i = 0 or z - i = 0 
==> z^2 = -i or z = i. 

In either case, we have |z| = 1, because 
(i) |z^2| = |-i| ==> |z|^2 = 1 ==> |z| = 1 (since |z| is non-negative real) 
(ii) |z| = |i| ==> |z| = 1.